OpenStudy (anonymous):

please help! Evaluate the following expression using the values given: Find 4x − y − 2z if x = −2, y = 3, and z = −2 −15< I thought it was this, but its not... 15 −7 7

OpenStudy (michaelalex292):

Input the values in the equation:\[4(-2) - (3) - 2(-2)\] And then solve

OpenStudy (anonymous):

I did.. I solved it, and I thought I solved it correctly... I even wrote the problem like that!

OpenStudy (michaelalex292):

Hmmm, let's see:\[(-8 ) - 3 = -11\] \[-11 + 4 = -7\] Remember that -2 times -2 is +4, so -7 is the answer

OpenStudy (anonymous):

oh thanks!

OpenStudy (michaelalex292):

No problem

OpenStudy (anonymous):

can you help with another one?

OpenStudy (michaelalex292):

Ok, but let me see it first

OpenStudy (michaelalex292):

What class are you taking btw?

OpenStudy (anonymous):

Evaluate the following: −3 − (−8). 5 −5 −11< wouldnt it be this? −18

OpenStudy (anonymous):

algebra 1

OpenStudy (michaelalex292):

Oh ok, so for this equation, this is what it looks like:\[-3 + -1(-8)\] When you have a negative sign right before the parentheses, it can be shown like above. Now can you solve it?

OpenStudy (anonymous):

Think of it like this \[\large (-3)+(-1(-1 \times 8)\] The two -1's cancel out leaving you with \[\large -3+8\]

OpenStudy (michaelalex292):

Yep, he's right

OpenStudy (anonymous):

so 5?

OpenStudy (michaelalex292):

Yes